OpenStudy (anonymous):

A periscope is 5 feet above the surface of the ocean. Through it can be seen a ship that rises to 50 feet above the water. To the nearest mile, the farthest away that the ship could be is ( ) miles

OpenStudy (anonymous):

5^2+50^2=25+2500=2525 \[\sqrt{2525}\] =505

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

505 was incorrect

OpenStudy (anonymous):

@helper99